OpenStudy (anonymous):

@Zale101 @ganeshie8 Which quadrant is point (12, –32) in? A. Quadrant IV B. Quadrant I C. Quadrant II D. Quadrant III

OpenStudy (anonymous):

@satellite73

jimthompson5910 (jim_thompson5910):

Hint: 

jimthompson5910 (jim_thompson5910):

the x axis is horizontal the y axis is vertical

OpenStudy (anonymous):

A

jimthompson5910 (jim_thompson5910):

if you have a positive x coordinate, then you're on the right half of the x axis 

jimthompson5910 (jim_thompson5910):

yep quadrant IV

OpenStudy (anonymous):

Which equation is shown on the graph? A. y = x – 2 B. y = x + 2 C. y = 2x D. y = 2x + 2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you forgot the graph